40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Dec 1 20:20:15 2016, gautam, Update, IMC, IMC loss measurement plan IMG_3471.JPG
    Reply  Fri Dec 2 16:40:29 2016, gautam, Update, IMC, 24V fuse pulled out 
       Reply  Mon Dec 5 01:58:16 2016, gautam, Update, IMC, IMC ringdowns IMCringdown.pdfIMCringdown_2.pdf
          Reply  Mon Dec 5 15:55:25 2016, gautam, Update, IMC, IMC ringdowns 7x
             Reply  Mon Dec 5 19:29:52 2016, gautam, Update, IMC, IMC ringdowns REFLanomaly.pdf
                Reply  Tue Dec 6 00:43:41 2016, gautam, Update, IMC, more IMC ringdowns 
                   Reply  Wed Dec 7 11:52:48 2016, ericq, Update, IMC, Partial IMC ringdowns IMCpartial.pngIMCpartial.zip
                      Reply  Thu Dec 8 19:01:21 2016, rana, Update, IMC, Partial IMC ringdowns 
                   Reply  Sun Dec 2 17:26:58 2018, gautam, Update, IMC, IMC ringdown fitting tauTheoretical.pdfringdownFit.pdfringdownScatter.pdfcavPole.pdf
                      Reply  Fri Dec 7 12:51:06 2018, gautam, Update, IMC, IMC ringdown fitting time_reflAndTrans.pdfcorner_transOnly.pdfcorner_reflAndTrans.pdf
Message ID: 12665     Entry time: Mon Dec 5 15:55:25 2016     In reply to: 12663     Reply to this: 12666
Author: gautam 
Type: Update 
Category: IMC 
Subject: IMC ringdowns 

As promised, here is the more detailed elog.


Part 1: AOM alignment and diffraction efficiency optimization

I started out by plugging in the input to the AOM driver back to the DS345 on the PSL table, after which I re-inserted the 24V fuse that was removed. I first wanted to optimize the AOM alignment and see how well we could cut the input power by driving the AOM. In order to investigate this, I closed the PMC, unlocked the PSL shutter, and dialed the PSL power down to ~100mW using the waveplate in front of the laser. Power before touching anything just before the AOM was 1.36W as measured with the Coherent power meter. 

The photodiode (PDA255) for this experiment was placed downstream of the 1%(?) transmissive optic that steers the beam into the PMC (this PD would also be used in Part 2, but has since been removed)...

Then I tuned the AOM alignment till I maximized the DC power on this newly installed PD. It would have been nicer to have the AOM installed on the mount such that the alignment screws were more easily accessible, but I opted against doing any major re-organization for the time being. Even after optimizing the AOM alignment, the diffraction efficiency was only ~15%, for 1V to the AOM driver input. So I decided to play with the AOM driver a bit.

Note that the AOM driver is powered by 24V DC, even though the spec sheet says it wants 28V. Also, the "ALC" input is left unconnected, which should be fine for our purposes. I opted to not mess with this for the time being - rather, I decided to tweak the RF adjust potentiometer on the front of the unit, which the spec sheet says can adjust the RF power between 1W and 2W. By iteratively tuning this pot and the AOM alignment, I was able to achieve a diffraction efficiency of ~87% (spec sheet tells us to expect 80%), in a switching time of ~130ns (spec sheet tells us to expect 200ns, but this is presumably a function of the beam size in the AOM). These numbers seemed reasonable to me, so I decided to push on. Note that I did not do a thorough check of the linearity of the AOM driver after touching the RF adjust potentiometer as Koji did - this would be relevant if we want to use the AOM as an ISS servo actuator, but for the ringdown, all that matters is the diffraction efficiency and switching time, which seemed satisfactory. 

At this point, I turned the PSL power back up (measured 1.36W just before the AOM). Before this, I estimated the PD would have ~10mW power incident on it, and I wanted it to be more like 1mW, so I I put an ND 1.0 filter on to avoid saturation.


Part 2: PMC "ringdown"

As mentioned in my earlier elog, we want the PMC to cut the light to the IMC in less than 1us. While I was at it, I decided to see if I could do a ringdown measurement for the PMC. For this, I placed two more PDs in addition to the one mentioned in Part 1. One monitored the transmitted intensity (PDA10CF, installed in the old 3f cancellation trial beam path, ~1mW incident on it when PMC is locked and well aligned). I also split off half the light to the PMC REFL CCD (2mW, so after splitting, PMC CCD gets 1mW through some ND filters, and my newly installed PD (PDA255) receives ~1mW). Unfortunately, the PMC ringdown attempts were not successful - the PMC remains locked even if we cut the incident light by 85%. I guess this isn't entirely surprising, given that we aren't completely extinguishing the input light - this document deals with this issue.... But the PMC transmitted intensity does fall in <200ns (see plot in earlier elog), which is what is critical for the IMC ringdown anyways. So I moved on.


Part 3: IMC ringdown

The PDA10CF installed in part 2 was left where it was. The reflected and transmitted light monitors were PDA255. The former was installed in front of the WFS2 QPD on the AS table (needed an ND1.0 filter to avoid damage if the IMC unlocks not as part of the ringdown, in which case ~6mW of power would be incident on this PD), while the latter was installed on the MC2 transmission table. We may have to remove the former, but I don't see any reason to remove the latter PD. I also ran a long cable from the MC2 trans table to the vertex area, which is where I am monitoring the various signals.

  

The triggering arrangement is shown below.

  

To actually do the ringdown, here is the set of steps I followed.

  1. Make sure settings on scope (X & Y scales, triggering) are optimized for data capture. All channels are set to 50ohm input impedance. The trigger comes from the "TTL" output of the DS345, whose "signal" output drives the AOM driver. Set the trigger to external, the mode should be "normal" and not "auto" (this keeps the data on the screen until the next trigger, allowing us to download the data via ethernet.
  2. The DS345 is set to output a low frequency (0.005Hz) square wave, with 1Vpp amplitude, 0.5V offset (so the AOM driver input is driven between 0V and 1V DC, which is what we want). This gives us ~100 seconds to re-lock the IMC, and download the data, all while chilling in the control room
  3. The autolocker was excellent yesterday, re-acquiring the IMC lock in ~30secs almost every time. But in the few instances it didn't work, turn the autolocker off (but make sure the MC2 tickle is on, it helps) and manually lock the IMC by twiddling the gain slider (basically manually do what the autolock script does). As mentioned above, you have ~100 secs to do this, if not just wait for 200secs and the next trigger...
  4. In the meantime, download the data (script details to follow). I've made a little wrapper script (/users/gautam/2016_12_IMCloss/grabChans.sh) which uses Tobin's original python script, which unfortunately only grabs data one channel at a time. The shell script just calls the function thrice, and needs two command line arguments, namely the base name for the files to which the data will be written, and an IP address for the scope...

It is possible to do ~15 ringdowns in an hour, provided the seismic activity is low and the IMC is in a good mood. Unfortunately, I messed up my data acquisiton yesterday, so I only have data from 2 ringdowns, which I will work on fitting and extracting a loss number from. The ringing in the REFL signal is also a mystery to me. I will try using another PDA255 and see if this persists. But anyways, I think we can exclude the later part of the REFL signal, and fit the early exponential decay, in the worst case. The ringdown signal plots have been uploaded to my previous elog. Also, the triggering arrangement can be optimized further, for example by using the binary output from one of our FEs to trigger the actual waveform instead of leaving it in this low frequency oscillation, but given our recent experience with the Binary Output cards, I thought this is unnecessary for the time being...

Data analysis to follow.


I have left all the PDs I put in for this measurement. If anyone needs to remove the one in front of WFS2, go ahead, but I think we can leave the one on the MC2 trans table there...

Attachment 2: AOMswitching.pdf  33 kB  | Hide | Hide all
AOMswitching.pdf
Attachment 6: electricalLayout.pdf  100 kB  Uploaded Mon Dec 5 17:03:26 2016  | Hide | Hide all
electricalLayout.pdf
ELOG V3.1.3-